I have a rectangular swimming pool that is 3 times as long as it is wide. Around the pool is a deck that is 2.5 meters wide and the area of the deck is 265 meters^2. How do I figure out the dimensions of the pool?
Answer by edjones(8007)

L=3w
(L+5)*(w+5)-(L*w)=265
(3w+5)*(w+5)-3w*w=265
3w^2+20w+25-3w^2=265
20w+25=265
20w=240
w=12
L=36
Ed